# An Enhanced Projection Pursuit Tree Classifier with Visual Methods for Assessing Algorithmic Improvements

### v1

@@ Abstract
This paper presents enhancements to the projection pursuit tree classifier and visual diagnostic methods for assessing their impact in high dimensions. The original algorithm uses linear combinations of variables in a tree structure where depth is constrained to be less than the number of classes—a limitation that proves too rigid for complex classification problems. Our extensions improve performance in multi-class settings with unequal variance-covariance structures and nonlinear class separations by allowing more splits and more flexible class groupings in the projection pursuit computation. Proposing algorithmic improvements is straightforward; demonstrating their actual utility is not. We therefore develop two visual diagnostic approaches to verify that the enhancements perform as intended. Using high-dimensional visualization techniques, we examine model fits on benchmark datasets to assess whether the algorithm behaves as theorized. An interactive web application enables users to explore the behavior of both the original and enhanced classifiers under controlled scenarios. The enhancements are implemented in the R package PPtreeExt.
@@ 1 Introduction
Loh (2014) provides an extensive overview of decades of research on classification and regression tree algorithms. Tree models are useful because they are simple, yet flexible, fast to compute, widely applicable and interpretable. A main drawback, though, most available methods use splits along single variables, which means that separations best made using a combination of variables are awkward for trees to fit. While there have been many suggestions it is a much harder problem to optimize on linear combinations, and thus there is less available successful software for oblique splits.
@@
One available method and software is the Projection Pursuit Tree (PPtree) algorithm (Y. D. Lee et al. 2013). It optimizes a projection pursuit index, for example the LDA (E.-K. Lee et al. 2005) or PDA (E.-K. Lee and Cook 2010) index among others, based on class information to identify one-dimensional projections that best separate the groups at each node. The PPtree tree structure is simpler than classic methods like rpart (Therneau and Atkinson 2025), as it restricts the depth to G1G-1 ($G$ is the number of classes). At each node, PPtree separates the data into two groups by first deciding which classes to group together; the data are split into left and right branches based on two group means. This produces a compact and interpretable tree implemented in PPtreeViz R package (E.-K. Lee 2018).
@@
One of the main advantages of PPtree is its ability to leverage the correlation between predictor variables to improve class separation. It has been shown to outperform even random forests (RF) in some scenarios. Projection pursuit addresses a known limitation of the original RF algorithm, in which oblique projections were allowed. While it sounds appealing, the algorithm is practically infeasible because it is based on randomly generated splits. Thus it is slow, and inefficient, for finding useful separating oblique splits.
@@
The compact nature of PPtree means that it fits simply, and does not require pruning. The resulting one-dimensional projections can be used to make visualizations of group separations. However, there are limitations that reduce its effectiveness in certain contexts. Due the the way classes are grouped, the prediction boundaries are often too close to one group, resulting in unnecessarily high error rates. This happens because grouping of classes likely produces unequal variance-covariances, making the pooled variance-covariance used by LDA inadequate. Moreover, when classes have non-elliptical variance-covariance, or are split into multiple clusters, the rigidity of the (G(G-1)-node structure, will not be sufficiently flexible.
A primary objective of this paper is to describe modifications to the PPtree algorithm that address these limitations. The algorithm is changed in two ways: (1) improving prediction boundaries by modifying the choice of split points-through class subsetting; and (2) increasing flexibility by allowing multiple splits per group. The modified algorithm is tested on a variety of benchmark data sets. These data sets include well-separated classes, the presence of outliers, and nonlinear boundaries. The performance is compared with the original PPtree algorithm and existing methods. Visual comparisons of decision boundaries, along with error rates, are used to further understand the model fits. The ultimate goal is to provide better building blocks for a projection pursuit-based classification random forest algorithm (da Silva, Cook, and Lee 2021) implemented in PPforest R package (da Silva, Cook, and Lee 2025).
@@
Proposing enhancements to algorithms is easy, and theoretically the modifications may seem sensible. In practice, they may not work as one envisions. Thus, an important part of developing algorithm modifications is providing validation that they are doing what is expected. This means going beyond predictive accuracy, to understanding the resulting models. A second objective of this paper is to provide visual diagnostics for the algorithm changes. This is two-fold: a shiny (Chang et al. 2025) app and tours to view the fits in high dimensions. The app is available in the PPtreeExt package (da Silva, Cook, and Lee 2026) along with the implementation of the modified algorithm. It allows users to generate 2D data sets with different types of separations and compare the model fits for the modified and the original algorithm. This was an essential tool in constructing the new algorithm, because it provides side-by-side visualization of prediction boundaries to identify data patterns in which the different methods perform better. The second visual diagnostic is to use tours available in the tourr package (Wickham et al. 2011) to examine the high-dimensional structure of data sets and the resulting model fits where the modified algorithm outperforms other classifiers. This provides an explanation for the performance comparison results, and intuition for why the modified algorithm performs better.
@@
The paper is organized as follows: Section 2 provides background on the original PPtree algorithm, including its structure and limitations. Section 3 describes the extensions to the algorithm. Section 4 presents the interactive visualization tool developed for exploring and comparing the algorithms. Section 5 presents the comparative analysis of algorithm performance, and visual diagnostics for the comparisons in Section 6. Finally, Section 7 discusses the changes, results of the comparison and potential future directions.
@@ 2 Background to PPtree
Figure 1 illustrates the original PPtree algorithm for $G=3$. Starting from the optimal projection, the green group-having the most distant mean—is relabeled as $g_{1}^{*}$, while the orange and violet groups are combined and relabeled as $g_{2}^{*}$. A second projection vector is then computed using all observations, following the same procedure applied to these two new super-groups. A split point is determined using one of the predefined rules, and observations are assigned to branches by comparing their projected values to the cutoff.
The algorithm computes a split point using one of several methods and the observations are allocated to groups based on whether their projected values fall below or above the cutoff. Since $g_{1}^{*}$ contains only one of the original classes, the left branch becomes a terminal node. Because $g_{2}^{*}$ still includes two original classes, the procedure is recursively repeated on the right branch.
Figure 1: Illustration of the original PPtree algorithm for $G=3$.
@@
Table 1: Rules for computing the split value, $c$, between two groups, on a data projection, using means, weighted mean, medians, standard deviations, interquartile ranges.
@@ 2.1 Algorithm
The PPtree algorithm (Y. D. Lee et al. 2013) follows a multi-step procedure to identify linear combinations of predictors that optimally separate classes in multi-class settings. Let $d_{n}=\{(\mathbf{x}i,y_{i})\}_{i=1}^{n}$ denote the dataset, where $\mathbf{x}_{i}$ is a $p$-dimensional vector of explanatory variables. The corresponding class label is $y_{i}\in\mathscr{G}$, with $\mathscr{G}=\{1,2,\ldots,G\}$, for $i=1,\ldots,n$. The construction of a PPtree classifier follows a four-step algorithm, described as follows:
PPtree Algorithm (4-step version)
@@
  • Step 1: Class Separation. Optimize a projection pursuit index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates all classes in the current node.
@@ 2.2 Less desirable aspects
The decision boundaries of a classifier reflect how the model partitions the feature space and adapt to class separations. Visualizing the decision boundaries provides valuable insights into model behavior and limitations. To illustrate the motivation for extending the PPtree algorithm, we present simple examples based on simulated datasets. Details on the simulation settings are provided in Section 4.
Figure 2 shows the decision boundaries produced by rpart (left) and PPtree (right). In this example, the class structure is best separated using linear combinations of the two variables. The rpart algorithm, which relies on axis-aligned splits, must approximate the ideal boundary using multiple steps. In contrast, PPtree generates an oblique split that more effectively captures the separation, although the resulting boundary between the orange and purple classes lies too close to one of them.
The first split isolates the violet class from the other two. A second partition then attempts to separate the green and orange groups. Although the group structure is similar across the three classes, the resulting decision boundaries are not parallel. This occurs because the initial split uses the combined information from the orange and green groups-treated as a super-group—to compute the projection and determine the splitting point.
Alternative rules for computing the split value (see Table LABEL:tbl-rules) have a substantial impact on the position and orientation of these decision boundaries.
Figure 2: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on two-dimensional simulated data. The boundaries generated by PPtree are oblique to the coordinate axes, capturing the linear association between the two variables.
Figure 3: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on two-dimensional simulated data. The orange class cannot be separated using a single linear partition, and PPtree fails to model it accurately because each original class must be assigned to a single terminal node.
Figure 3 displays a simulated example that highlights the limited flexibility of the decision boundaries produced by PPtree. In this scenario, the orange class cannot be separated using a single linear partition, and PPtree fails to model it accurately because each original class must be assigned to a single terminal node. In contrast, the more traditional rpart algorithm performs better in this case, as its recursive structure allows it to accommodate the relatively simple non-linear separation.
@@ 3 PPtree extensions
The algorithm has been extended in two main ways: by modifying the rules used to determine decision boundaries and by allowing multiple splits per group. Two new approaches are introduced for computing the optimal split value. To support multiple splits within a group, additional stopping rules must be incorporated into the algorithm.
@@ 3.1 Modification 1: Subsetting classes to produce better boundaries
The first modification targets the third step of the original algorithm. Rather than combining all remaining classes into a single super-class, only the two closest ones—identified by the smallest difference in their projected means—are used to compute the next projection and define the split point
$3^{*}$ Find an optimal one-dimensional projection $\alpha^{**}$, using a subset of $\{(\mathbf{x_{i}},y_{i}^{*})\}_{i=1}^{n}$ to separate the two class problem $g_{1}^{*}$ and $g_{2}^{*}$. To this second projection only information from the two closest groups, one from $g_{1}^{*}$ and the other from $g_{2}^{*}$ are used. The best separation of $g_{1}^{*}$ and $g_{2}^{*}$ is determined by $\boldsymbol{\alpha}^{*T}\mathbf{\bar{x}}_{g1^{*}}<c$; then $g_{1}^{*}$ is assigned to the left node else assign $g_{2}^{*}$ to the right node, where $\mathbf{\bar{x}}_{g1^{*}}$ is the mean of $g_{1}^{*}$.
Figure 4 illustrates the modified version of the algorithm for a three-class classification problem. After identifying the initial optimal projection, instead of combining two groups into a single super-group, the group furthest from the others is temporarily excluded. In this example, class $g_{3}$ is dropped from the current node. A new projection is then computed using only classes $g_{1}$ and $g_{2}$ (relabeled as $g_{1}^{*}$ and $g_{2}^{*}$ ), and one of the eight available rules is applied to determine the optimal split point. Subsequently, class $g_{3}$ is reintroduced, and a new projection is computed to best separate it from $g_{2}$, completing the tree construction.
Figure 6 shows the effect of this modification on the decision boundaries obtained from the simulated dataset presented in Figure 2.
Figure 4: Illustration of the algorithm with Modification 1 applied to a three-class problem. At each node, the second one-dimensional projection is computed using only the two closest groups to determine the best projection direction and split point.
@@ 3.2 Modification 2: Multiple splits per class based on entropy reduction
This modification introduces a new approach for selecting the split value, based on the impurity of the resulting groups. This approach resembles the impurity-based criteria used in the rpart algorithm, but it operates on projected data rather than individual features. Currently, the only implemented impurity measure is negative entropy:
@@
where $p_{js}$ is the proportion of observations from class $j$ in subset $s$, and $G$ is the number of classes. Higher values of $E(s)$ indicate greater class impurity. When all observations in the subset belong to the same class, the entropy reaches its minimum value of $E(s)=0$, indicating a pure node. To evaluate the quality of a candidate split, the impurity values of the resulting left and right subsets are combined using a weighted average of the entropie:
@@
where $n_{L},n_{R}$ are the number of observations in the left and right subsets, respectively. The combined impurity, $E(s_{L},s_{R})$ is computed for all possible split points-defined as the midpoints between consecutive ordered projected values-denoted $\mathscr{C}=\{c_{1},...,c_{n_{L}+n_{R}-1}\}$. The split point that minimizes the combined entropy is selected:
@@
This represents a major departure from the original PPtree structure, aiming to enhance flexibility in modeling non-linear class boundaries by allowing multiple splits within a class. Instead of relying on summary statistics at the class level, this approach shifts the focus to local differences between individual observations. To support this flexibility, additional control mechanisms are required, particularly new stopping criteria. While this is conceptually similar to the rpart algorithm, a key distinction is that larger sample sizes per node are needed to reliably estimate projection coefficients.
Figure 5 illustrates this modified version of the algorithm, which follows the steps outlined below.
PPtree Algorithm with Entropy-Based Splitting
@@
  • Step 1: Class Separation. Optimize a projection pursuit index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates the classes in the current node.
@@
4 Algorithm comparison

### v2

@@ Abstract
This paper presents enhancements to the projection pursuit tree classifier and visual diagnostic methods for assessing their impact in high dimensions. The original algorithm uses linear combinations of variables in a tree structure where depth is constrained to be less than the number of classes—a limitation that proves too rigid for complex classification problems. Our extensions improve performance in multi-class settings with unequal variance-covariance structures and nonlinear class separations by allowing more splits and more flexible class groupings in the projection pursuit computation. Proposing algorithmic improvements is straightforward; demonstrating their actual utility is not. We therefore develop two visual diagnostic approaches to verify that the enhancements perform as intended. Using high-dimensional visualization techniques, we examine model fits on benchmark datasets to assess whether the algorithm behaves as theorized. An interactive web application enables users to explore the behavior of both the original and enhanced classifiers under controlled scenarios. The enhancements are implemented in the R package PPtreeExt, which is available on CRAN.
Natalia da Silva1   Dianne Cook2   Eun-Kyung Lee3
1Instituto de Estadística (IESTA), Universidad de la República, Montevideo, Uruguay
2Department of Econometrics and Business Statistics, Monash University, Melbourne, Australia
3Department of Statistics, Ewha Womans University, Seoul, Republic of Korea
Keywords: classification, heterogeneous variance, machine learning, projection pursuit, separability, tree classifier
@@ 1 Introduction
Loh (2014) provides an extensive overview of decades of research on classification and regression tree algorithms. Tree models are useful because they are simple, yet flexible, fast to compute, widely applicable, and interpretable. A main drawback, though, is that most available methods use splits along single variables, making separations on a combination of variables harder to capture. The reason is that it is a much harder problem to optimize the search for linear combinations, and thus, there is a lack of useful software that provides oblique splits.
@@
One available method and software is the projection pursuit tree (PPtree) (Lee et al. 2013). It optimizes a projection pursuit (PP) index, for example, the LDA (Lee et al. 2005) or PDA (Lee and Cook 2010) index among others, based on class information to identify one-dimensional projections that best separate the groups at each node. The PPtree tree structure is simpler than classic methods like rpart (Therneau and Atkinson 2025), as it restricts the tree depth to G1G-1 ($G$ is the number of classes). At each node, PPtree separates the data into two groups by first deciding which classes to group together; the data are split into left and right branches based on two group means. This produces a compact and interpretable tree implemented in the PPtreeViz R package (Lee 2018).
@@
One of the main advantages of PPtree is its ability to leverage the correlation between predictor variables to improve class separation. In some scenarios, it has been shown to outperform other classifiers, including random forests (RF) (Breiman 2001). PP addresses a known limitation of the original RF algorithm. It included oblique projections, but because it is based on randomly generated splits was practically infeasible. It is an inefficient algorithm for finding useful oblique splits.
@@
The compact nature of PPtree means that it fits simply and does not require pruning. The resulting one-dimensional projections can be used to make visualizations of group separations. However, there are limitations that reduce its effectiveness. In multiclass problems, due to the way classes are grouped, the prediction boundaries are often too close to one group, resulting in unnecessarily high error rates. This happens because grouping of classes likely produces unequal variance-covariances, making the pooled variance-covariance used by LDA inadequate. Secondly, when classes have non-elliptical variance-covariance, or have multiple clusters per class, the rigidity of the (G(G-1)-node structure is not sufficiently flexible. The algorithm is modified to address these limitations in two ways: (1) improving prediction boundaries by modifying the choice of split points through class subsetting; and (2) increasing flexibility by allowing multiple splits per group.
@@
Proposing enhancements to algorithms is easy, and theoretically, the modifications may seem sensible. In practice, they may not work as one envisions. Thus, an important part of developing algorithm modifications is providing validation that they are doing what is expected, or not. This means going beyond predictive accuracy to understanding the resulting models. A second objective of this paper is to provide visual diagnostics for the algorithm changes. This is two-fold: a shiny (Chang et al. 2025) app and tours to view the fits in high dimensions. The app is available in the PPtreeExt package (da Silva et al. 2026) along with the implementation of the modified algorithm. It allows users to generate 2D data sets with different types of separations and compare the model fits for the modified and the original algorithm. This was an essential tool in constructing the new algorithm, because it provides a side-by-side visualization of prediction boundaries to identify data patterns in which the different methods perform better. The second visual diagnostic is to use tours available in the tourr package (Wickham et al. 2011) to examine the high-dimensional structure of data sets and the resulting model fits, where the modified algorithm outperforms other classifiers. This explains the performance comparison results and intuition for why the modified algorithm performs better.
@@
The paper is organized as follows: Section 2 provides background on the original PPtree algorithm, including its structure and limitations. Section 3 describes the extensions to the algorithm. Section 4 contains the examples on visually and numerically digesting the enhanced algorithm’s performance. Finally, Section 5 discusses the changes, results of the comparison, and potential future directions.
@@ 2 Background to PPtree
Figure 1 illustrates the original PPtree algorithm for $G=3$. Starting from the optimal projection, the left-side, most distant, cluster (green circles) is relabeled as $g_{1}^{*}$, while the other two clusters are combined and relabeled as $g_{2}^{*}$. A second projection vector is then computed using all observations, following the same procedure applied to these two new super-groups. A split point is determined using one of the predefined rules, and observations are assigned to branches by comparing their projected values to the cutoff.
The algorithm computes a split point using one of several methods, and the observations are allocated to groups based on whether their projected values fall below or above the cutoff. Since $g_{1}^{*}$ contains only one of the original classes, the left branch becomes a terminal node. Because $g_{2}^{*}$ still includes two original classes, the procedure is recursively repeated on the right branch.
Figure 1: Illustration of the original PPtree algorithm for $G=3$. In step 1, $P$ optimization decides on the best linear projection. Step 2 combines two groups to make a two-group classification. The $P$ optimization is computed for two group problem, and an appropriate split is decided. This is repeated for the remaining two groups to get the final model containing two splits. The locations where modifications will be made are circled.
@@
Table 1: Rules for computing the split value, $c$, between two groups, on a data projection, using means, weighted mean, medians, standard deviations, and interquartile ranges.
@@ 2.1 Algorithm
The PPtree algorithm (Lee et al. 2013) follows a multi-step procedure to identify linear combinations of predictors that optimally separate classes in multi-class settings. Let $d_{n}=\{(\mathbf{x}i,y_{i})\}_{i=1}^{n}$ denote the dataset, where $\mathbf{x}_{i}$ is a $p$-dimensional vector of explanatory variables. The corresponding class label is $y_{i}\in\mathscr{G}$, with $\mathscr{G}=\{1,2,\ldots,G\}$, for $i=1,\ldots,n$. The construction of a PPtree classifier follows a four-step algorithm, as follows:
@@
  • Step 1: Class Separation. Optimize a PP index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates all classes in the current node.
@@ 2.2 Less desirable aspects
The decision boundaries of a classifier reflect how the model partitions the feature space and adapts to class separations. Visualizing the decision boundaries provides valuable insights into model behavior and limitations. To illustrate the PPtree algorithm relative to the classical rpart and the motivation for improving PPtree, we use two simulated three-class 2D data sets. Data set LC has separations on linear combinations of variables (Figure 2), and data set MC has multiple clusters in a class (Figure 3). The three classes are indicated by squares (S, violet), circles (C, green), and triangles (T, orange).
Figure 2 shows the decision boundaries produced by rpart (left) and PPtree (right), and percentage error, for the LC data. The class clusters are best separated using linear combinations of the two variables. The rpart algorithm, which relies on axis-aligned splits, must approximate the ideal boundary using rectangular steps. Even with many steps, it cannot be as efficient as the oblique splits provided by PPtree. The two steps of PPtree are to split the S cluster (violet) first, and then split C (green) and T (orange). Although the group structure is similar across the three classes, the resulting decision boundaries are not parallel. This occurs because the initial split uses the combined information from the C and T clusters - treated as a super class - to compute the projection and determine the splitting point. This also results in the boundary between T and S being too close to T. Alternative rules for computing the split value (see Table LABEL:tbl-rules) can impact the position and orientation of these decision boundaries, but in this case does not fix the deficiency.
Figure 2: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on the LC data. The boundaries generated by PPtree are oblique to the coordinate axes, appropriately capturing the separations as linear combinations of the two variables. Even with many splits, the rpart algorithm cannot elegantly classify this data.
Figure 3: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on the MC data. The T class (orange triangles) cannot be separated using a single linear partition, and PPtree fails to model it accurately because each class must be assigned to a single terminal node. This illustrates the need for allowing more splits in PPtree to capture unusually shaped clusters and separations involving combinations of variables.
Figure 3 compares the rpart and PPtree decision boundaries for the data set MC, where class T has two separated clusters. It highlights the limited flexibility of the decision boundaries produced by PPtree. In this scenario, the T cannot be separated using a single linear partition, and PPtree fails to capture it because the tree depth is limited to a single terminal node for each class. In contrast, the more traditional rpart algorithm performs better in this case, as its recursive structure allows it to accommodate the relatively simple non-linear separation. However, rpart fails to capture the separation between classes as linear combinations. This shows why more splits should be allowed for the PPtree algorithm.
@@ 3 PPtree extensions
The algorithm has been extended in two main ways, as indicated by the annotation in Figure 1. Section 4.1 explains how we can compare the changes these modifications induce on the boundaries relative to rpart and PPtree using the Shiny app.
@@ 3.1 Modification 1: Subsetting classes to produce better boundaries
The first modification targets the re-grouping step of the original algorithm. Rather than combining all remaining classes into a single super-class, only the two closest ones, as identified by the smallest difference in their projected means, are used to compute the next projection and define the split point. This step is now as follows:
  • Step 2: Binary Relabeling. Find an optimal one-dimensional projection $\alpha^{**}$, using a subset of $\{(\mathbf{x_{i}},y_{i}^{*})\}_{i=1}^{n}$ to separate the two class problem $g_{1}^{*}$ and $g_{2}^{*}$. To this second projection only information from the two closest groups, one from $g_{1}^{*}$ and the other from $g_{2}^{*}$ are used. The best separation of $g_{1}^{*}$ and $g_{2}^{*}$ is determined by $\boldsymbol{\alpha}^{*T}\mathbf{\bar{x}}_{g1^{*}}<c$; then $g_{1}^{*}$ is assigned to the left node else assign $g_{2}^{*}$ to the right node, where $\mathbf{\bar{x}}_{g1^{*}}$ is the mean of $g_{1}^{*}$.
In Figure 1 this would correspond to dropping class $g_{3}$ from the current node. The projection is then computed using only classes $g_{1}$ and $g_{2}$ (relabeled as $g_{1}^{*}$ and $g_{2}^{*}$ ). The choice of the existing eight available rules is then applied to determine the optimal split point, to create the terminal node for $g_{1}$. Subsequently, class $g_{3}$ is reintroduced, and a new projection is computed to best separate it from $g_{2}$. The split is made on this set of observations, creating two more terminal nodes, completing the tree construction.
@@ 3.2 Modification 2: Multiple splits per class based on entropy reduction
To allow for multiple splits, various changes are needed: the split is made on the optimal projection, not after groups are combined, and new stopping rules are defined to prevent over-fitting. Figure 4 illustrates the change to the algorithm for this modification.
In PPtree, the decision boundary split is made using the means for the two groups, just like is done with linear discriminant analysis. This modification changes this to follow the rpart algorithm, minimizing entropy, with the difference being that it is done on the optimal 1-D projected data. That is, the modification computes the negative entropy criteria:
@@
where $p_{js}$ is the proportion of observations from class $j$ in subset $s$, and $G$ is the number of classes. Lower values of $E(s)$ indicate purity, reaching zero when all observations in the subset belong to the same class. The impurity value for a split is calculated by combining the values of the left and right subsets:
@@
where $n_{L},n_{R}$ are the number of observations in the left and right subsets, respectively. Across all possible splits, the lowest combined impurity, $E(s_{L},s_{R})$, is computed for all possible split points, defined as the midpoints between consecutive ordered projected values, denoted $\mathscr{C}=\{c_{1},...,c_{n_{L}+n_{R}-1}\}$. The split point that minimizes the combined entropy is selected:
@@
In addition, three stopping rules are introduced, which are similar to rpart: the node is pure, there are fewer than a specified number of observations in the node, or the reduction in impurity is less than a specified tolerance value. These changes provide the modified algorithm to be:
@@
  • Step 1: Class Separation. Optimize a PP index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates the classes in the current node.
@@
4 Comparative performance
Three approaches are used to examine the effectiveness of the modifications. The first is a Shiny app which generates structures in 2-D and shows the model fits for three algorithms: rpart, PPtree, and the PPtree extensions. It was developed and used to decide on effective changes to the algorithm. The second is a simulation study on 94 benchmark data sets to compare test errors. Lastly, to understand where, why, and how the extensions work, we examine the data visually in high dimensions along with the fitted models.
4.1 Exploring the algorithm changes interactively